# Preconditioning without a preconditioner using randomized block Krylov subspace methods

### v2

@@ B.4 What if CG uses some reorthogonalization?
Comparing vanilla CG with Nyström PCG as in Fig. 10(a) is perhaps a bit unfair to CG; Nyström PCG incurs costs in building, storing, and applying the preconditioner. To put the algorithms on more equal footing, we might allow CG to use a similar amount of storage and arithmetic operations to do some form of reorthogonalization. While there are many complicated schemes based on detailed analyses of the Lanczos algorithm in finite precision arithmetic [51, 58], a simple scheme is to perform re-orthogonlization only for the first $s\ell$ iterations, and then to continue orthogonalize against these vectors in subsequent iterations. In this case the cost of reorthogonalization during the first $s\ell$ iterations will be $O(ds^{2}\ell^{2})$ operations (the same as the cost to build the the Nyström preconditioner) and subsequently orthogonalizing against these vectors will require $O(ds\ell)$ operations per iteration, the same as the cost to apply the Nyström preconditioner to a vector.
In Fig. 10(b) we show the behavior of CG with full-reorthogonalization for the first 30 iterations and then orthogonalizing subsequent Krylov basis vectors against these first thirty vectors. Compared to Fig. 10(a), we observe that CG now converges as rapidly as Nyström PCG.
@@
In Fig. 2 we compare the performance of these methods when we do not use full-reorthogonalzation. In particular, we run Nyström PCG with $s=3$ without any reorthogonalization. Building the preconditioner requires orthogonalizing $s\ell$ vectors. For block-CG and CG use full reorthogonalization until we have a set of $3\ell$ vectors and then continue to orthogonalize only against these vectors. For reference, we also display the convergence of block-CG and CG if no reorthogonalization is used at all. Unfortunately, far less is known about the behavior of block-CG and block-Lanczos in finite precision arithmetic than is known about their single-vector counterparts.
Figure 11: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-loads for block-CG (Refer to caption) with reorthogonalization for 3 iterations, CG (Refer to caption) with reorthogonalization for $3\ell$ iterations, and Nyström PCG with $s=3$ (Refer to caption) without any reorthgonalization. Light curves show convergence of block-CG and CG with no reorthgonalization. The test problems are the same as Fig. 2.
@@ Appendix C Omitted proofs
(54) $\mathbf{E}=\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle,\quad\widehat{\mathbf{A}}_{\mu}=\mathbf{A}\langle\mathbf{K}\rangle+\mu\mathbf{I}.$
@@
(55) $\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$
@@
(56) $\widehat{\mathbf{A}}_{\mu}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}+\mu\mathbf{I}=\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$
@@
(57) $\|\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|\leq\|\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|+\|\mathbf{P}_{\mu}^{-1/2}\mathbf{E}\mathbf{P}_{\mu}^{-1/2}\|.$
@@
(58) $\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}=(\theta+\mu)\mathbf{U}\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$
@@
We now derive a lower bound for the eigevnalues of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$. Observe that
(60) $\hskip-30.00005pt\|(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})^{-1}\|$
(61) $=\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|$
(62) $=\Big{\|}\mathbf{A}_{\mu}^{-1/2}\Big{(}\frac{1}{\theta+\mu}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\Big{)}\mathbf{A}_{\mu}^{-1/2}\Big{\|}$
(63) $\leq\frac{1}{\theta+\mu}\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|+\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|.$
@@
(64) $\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|\leq 1.$
@@
(65) $\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|\leq\|\mathbf{A}_{\mu}^{-1}\|=\frac{1}{\lambda_{n}+\mu}.$
(66) $\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|\leq\frac{1}{\theta+\mu}+\frac{1}{\lambda_{n}+\mu}.$
@@ C.2 Bound on Elliptic integral
(69) $K(m)=\frac{1}{\sqrt{1-m}}K\left(\frac{-m}{1-m}\right),$
@@
(70) $K(1-x)\leq\frac{1}{\sqrt{x}}K\left(\frac{x-1}{x}\right).$
@@
(71) $\sqrt{x}K(1-x)\leq K\left(1-\frac{1}{x}\right),$
and if $x>1$ then $1-1/x\in(0,1)$ so we can apply Eq. 68 to obtain the bound
(72) $\sqrt{x}K(1-x)\leq\log\left(\frac{4}{\sqrt{1/x}}\right)\left(1+\frac{1/x}{4}\right).$

### v3

@@ B.4 What if CG uses some reorthogonalization?
Comparing vanilla CG with Nyström PCG as in ˜10(a) is perhaps a bit unfair to CG; Nyström PCG incurs costs in building, storing, and applying the preconditioner. To put the algorithms on more equal footing, we might allow CG to use a similar amount of storage and arithmetic operations to do some form of reorthogonalization. While there are many complicated schemes based on detailed analyses of the Lanczos algorithm in finite precision arithmetic [57, 65], a simple scheme is to perform re-orthogonlization only for the first $s\ell$ iterations, and then to continue orthogonalize against these vectors in subsequent iterations. In this case the cost of reorthogonalization during the first $s\ell$ iterations will be $O(ds^{2}\ell^{2})$ operations (the same as the cost to build the the Nyström preconditioner) and subsequently orthogonalizing against these vectors will require $O(ds\ell)$ operations per iteration, the same as the cost to apply the Nyström preconditioner to a vector. In ˜10(b) we show the behavior of CG with full-reorthogonalization for the first 22 iterations and then orthogonalizing subsequent Krylov basis vectors against these first thirty vectors. Compared to ˜10(a), we observe that CG now converges as rapidly as Nyström PCG.
@@
In ˜2 we compare the performance of these methods when we do not use full-reorthogonalzation. In particular, we run Nyström PCG with $s=3$ without any reorthogonalization. Building the preconditioner requires orthogonalizing $s\ell$ vectors. For block-CG and CG use full reorthogonalization until we have a set of $3\ell$ vectors and then continue to orthogonalize only against these vectors. For reference, we also display the convergence of block-CG and CG if no reorthogonalization is used at all. Unfortunately, far less is known about the behavior of block-CG and block-Lanczos in finite precision arithmetic than is known about their single-vector counterparts.
Figure 11: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-loads for block-CG (Refer to caption) with reorthogonalization for 3 iterations, CG (Refer to caption) with reorthogonalization for $3\ell$ iterations, and Nyström PCG with $s=3$ (Refer to caption) without any reorthgonalization. Light curves show convergence of block-CG and CG with no reorthgonalization. The test problems are the same as ˜2.
@@ Appendix C Omitted proofs
$\mathbf{E}=\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle,\quad\widehat{\mathbf{A}}_{\mu}=\mathbf{A}\langle\mathbf{K}\rangle+\mu\mathbf{I}.$
@@
$\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$
@@
$\widehat{\mathbf{A}}_{\mu}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}+\mu\mathbf{I}=\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$
@@
$\|\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|\leq\|\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|+\|\mathbf{P}_{\mu}^{-1/2}\mathbf{E}\mathbf{P}_{\mu}^{-1/2}\|.$
@@
$\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}=(\theta+\mu)\mathbf{U}\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$
@@
We now derive a lower bound for the eigevnalues of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$. Using that $(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})^{-1}=\mathbf{P}_{\mu}^{1/2}\mathbf{A}_{\mu}^{-1}\mathbf{P}_{\mu}^{1/2}$ is similar to $\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}$, the definition of $\mathbf{P}_{\mu}$, and the triangle inequality,
$\hskip-30.00005pt\|(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})^{-1}\|$
$=\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|$
$=\Big\|\mathbf{A}_{\mu}^{-1/2}\Big(\frac{1}{\theta+\mu}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\Big)\mathbf{A}_{\mu}^{-1/2}\Big\|$
$\leq\frac{1}{\theta+\mu}\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|+\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|.$
@@
$\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|\leq 1.$
@@
$\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|$ $\leq\|\mathbf{A}_{\mu}^{-1}\|=\frac{1}{\lambda_{n}+\mu}$
$\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|$ $\leq\frac{1}{\theta+\mu}+\frac{1}{\lambda_{n}+\mu}.$
@@ C.2 Bound on Elliptic integral
$K(m)=\frac{1}{\sqrt{1-m}}K\left(\frac{-m}{1-m}\right),$
@@
$K(1-x)\leq\frac{1}{\sqrt{x}}K\left(\frac{x-1}{x}\right).$
@@
$\sqrt{x}K(1-x)\leq K\left(1-\frac{1}{x}\right),$
and if $x>1$ then $1-1/x\in(0,1)$ so we can apply ˜21 to obtain the bound
$\sqrt{x}K(1-x)\leq\log\left(\frac{4}{\sqrt{1/x}}\right)\left(1+\frac{1/x}{4}\right).$